OpenStudy (theraggedydoctor):

Simplify each expression. a. sqrt(7x^3) * sqrt(14x) b. (64)^-2/3

OpenStudy (anonymous):

\[64^{-\frac{2}{3}}=\frac{1}{\sqrt[3]{64}}\] is a start

OpenStudy (anonymous):

actually \[64^{-\frac{2}{3}}=\frac{1}{\sqrt[3]{64}^2}\]

OpenStudy (anonymous):

the cube root of 64 is 4, 4 squared is 16 finishes it off

OpenStudy (jhannybean):

And in the a. you can combine everything under one square root function.\[\sqrt{7x^3}\cdot \sqrt{14x} = \sqrt{ (7 \cdot 14)x^4}\]

OpenStudy (triciaal):

the post for a. can be simplified